[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Fri Jul 21, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Fri Jul 21 09:12:47 UTC 2017


HSU Shift Log for Fri Jul 21 2017 05:12:40 GMT-0400 (EDT)

Shift: A

Forecaster: Other

Technical Issues Summary:

ATCF issues: None

N-AWIPS issues: Rebooted Tuna and Sawfish. Tuna screen turned green before  
reboot (it seemed that the video driver replaced anything black with  
green), but no problems after reboot.

AWIPS-2/NCP issues: None

Other Techincal issues: None

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170721/25303148/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list